# A Watermark-Conditioned Diffusion Model for IP Protection

### v1

@@ 4.2 Rooting Fingerprints in Diffusion Process
End-to-end Fine-tuning
Instead of fine-tuning the entire diffusion model, we keep the weights of the remaining model unchanged and only fine-tune the first input block. We adopt this approach for two reasons. First, we observe that fine-tuning the first block is sufficient for effective watermark injection. Our method achieves comparable performance to fine-tuning the entire architecture but with faster speed and lower memory cost. Moreover, we have empirically observed that fine-tuning the entire model can lead to unstable generation, where the watermarked content deviates from the original structure of the generated content. We provide a detailed analysis of fine-tuning various model subsections in the Appendix. Formally, we separate $\theta$ into the parameters of the first input block as $\theta_{head}$ and the remaining weights $\theta_{tail}$, i.e. $\theta=\{\theta_{head},\theta_{tail}\}$. During fine-tuning, only $\theta_{head}$ is optimized while $\theta_{tail}$ is fixed. To sum up, we incorporate both optimization objectives above and formulate the fine-tuning process as:
$\min_{\theta_{head}}\mathbb{I}(t\leq\tau)(\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{i}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t}))+\eta\mathcal{L}_{m}(\mathcal{D}(\hat{\mathbf{x}}_{0}^{t}),w_{i}))$ (7)
$+\mathbb{I}(t>\tau)\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{null}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t})),$
@@
Figure 3: Overview of our proposed watermarking framework. The top two figures illustrate our fine-tuning process. For $t>\tau$, we solely focus on preserving image consistency. To achieve this, we incorporate a commonly shared null watermark and align the generated output with the original diffusion model. For $t\leq\tau$, we integrate the normal watermark and introduce an additional message retrieval loss to embed watermarking into the generation process. The inference stage is depicted below, where we inject the null watermark when $t>\tau$ and transition it to the payload watermark when $t\leq\tau$.
@@ 4.3 Detection and Identification
Once the diffusion model is fingerprinted, the generated content for each $u_{i}$ will contain specific information that is conditioned by $w_{i}$. We can then recover the source watermark $w_{s}=\mathcal{D}(\mathbf{p})$ using the pre-trained $\mathcal{D}$ model. Once we have recovered $w_{s}$, we can utilize it for both detection and identification. To detect whether $\mathbf{p}$ belongs to our model, we set the bit threshold $\tau_{b}$ and calculate the number of matched bits $M_{i}=w_{s}\oplus w_{i}$, where $\oplus$ represents the XOR function. If $M_{i}>\tau_{b}$, we will conclude that $\mathbf{p}$ is generated from our model; otherwise, it is generated from other sources. For identification among $m$ users, we can determine the owner by finding the corresponding user ID whose watermark $w_{i}$ has the best match with $w_{s}$, which can be formulated as follows:
$\arg\min_{i}M_{i},\ \ i\in\{1\dots m\}.$ (8)
@@ 5 Experiments
To evaluate the identification performance of our proposed WaDiff, we implement our method on two widely used diffusion models: a text-to-image latent diffusion model and an unconditional diffusion model. We conduct comprehensive comparisons with other watermarking strategies including a traditional post-hoc strategy DwtDct [5], a deep-learning-based steganography technique StegaStamp [30], which we employed in our pre-training stage, and a recently proposed training-free framework Tree-Ring [32]. Note that we have not included the Stable Signature [10] in our comparison since it considered the model distribution setting which is different from ours. Besides, it necessitates re-training diverse latent decoders, which is not scalable in our identification task. We also conduct thorough ablation studies and provide a detailed analysis in the following sections.
@@
5.1 Experimental Settings
Model and Dataset
This section introduces the model and dataset utilized in our experiments. For the text-to-image latent diffusion model, we utilize the Stable Diffusion V1.4 as our pre-trained model [^1] https://github.com/CompVis/stable-diffusion, and for the unconditional diffusion model, we utilize the 256 $\times$ 256 ImageNet diffusion model [^2] https://github.com/openai/guided-diffusion. To fine-tune the diffusion models, we randomly select 5000 images with corresponding text descriptions from the training set of MS-COCO 2014 [17] for the Stable Diffusion and 5000 images from the training set of ImageNet [7] for the 256x256 ImageNet diffusion model.
Trace $\text{Trace Acc}=\frac{N_{\text{correct}}}{N_{\text{total}}},$ (9)
Implementation Details
All experiments were conducted utilizing 8 NVIDIA A100 GPUs. The fine-tuning of both diffusion models involved employing an AdamW optimizer [19] with a fixed learning rate of $1e^{-4}$. For the Stable Diffusion model, we set $\tau$ to 500 and $\eta$ to 0.05, while for the ImageNet diffusion model, we selected $\tau$ as 400 and $\eta$ as 0.25. Notably, in addition to directly aligning the predicted noise of the ImageNet diffusion model with the original model, we empirically found that aligning the single-step-reverse images resulted in improved image quality. Implementation details and additional hyperparameter evaluations are deferred in the Appendix. We fine-tune 40 epochs for the Stable Diffusion model and 25 epochs for the ImageNet diffusion model. For the Stable Diffusion model, we set the default guidance scale to 7.5 and used text descriptions from the validation set of MS-COCO 2014 [17] as prompts. We adopt the watermark length to 48 by default, which is commonly used in previous work [10], and provide experiments of other lengths in the Appendix. We adopted the DDIM [29] sampler with 50 sampling steps as default for both models. More implementation details on training baseline methods and the watermark decoder pre-training are in the Appendix.
@@ 5.2 Detection and Identification Results
In this section, we present the main experimental results of our method. For the detection task, we use 5000 watermarked images along with another 5000 clean images sampled from the original pre-trained diffusion model to calculate the AUC. As for owner identification, we evaluate our method using different sizes of user pools, ranging from ten thousand to one million users. For each user pool, we randomly select 1000 users and generate 5 images per user, resulting in a total of 5000 images. The tracing accuracy is then calculated based on these watermarked images. For SSIM calculation, we first randomly select 200 distinct initial noises and generate a group of 5 images for each noisy vector with different keys. Within each group, the SSIM metric was computed by comparing the similarity between each pair of images. To calculate the FID difference, we evaluate the generative images (for both watermarked and originally generative contents) on the MS-COCO 2014 training set for Stable Diffusion and on the ImageNet [7] training set for the ImageNet diffusion model. We present the experimental results in Table 1. The results demonstrate that our method achieves comparable performance in both detection and identification tasks when compared to the post-hoc StegaStamp, achieving a tracing accuracy of 97.71% and an AUC of 1 on average. We defer further discussion on the comparison with post-hoc methods to the Appendix. In terms of image consistency, our method achieves an average of 0.998 SSIM between pairs of images with different watermarks, which significantly surpasses that of Tree-Ring. Besides, our method only imposes a negligible impact on the original generation quality by a slight increase on the original FID, which is comparable to the post-hoc watermarks. It is worth noting that the Tree-Ring achieves nearly zero performance on the identification task. This might be attributed to the imprecise latent inversion and the continuous watermarking space, which make watermarks less distinguishable when multiple users are involved.
@@ 5.3 Watermark Robustness Analysis
To evaluate our watermarking strategy against potential data augmentations, we adopt five commonly used augmentations including image resizing, image blurring, color jitter, Gaussian noising, and JPEG compression. Specifically, for resizing, we randomly resize the width and height of images to a range of 30% to 80% of their original size; for image blurring, we adopt the kernel size to 20; for color jitter, we randomly select from saturation (factor 1.5), contrast (factor 1.5) and sharpness (factor 1.5); for Gaussian noising, we add a Gaussian noise with $\sigma=0.1$; for the JPEG compression, we select the compression quality as 50. We also consider a combinational augmentation that incorporates image resizing, JPEG compression, and color jitter simultaneously. We report the identification performance under various augmentations in Table 2. The experimental results demonstrate the overall robustness of our method against diverse image augmentations, yielding an average tracing accuracy of 88.05% for the Stable Diffusion model and 89.7% for the ImageNet diffusion model. It is noteworthy that our method maintains an average AUC of 0.999 under various data augmentations and achieves identification performance of over 75% among one million users even under combined data augmentation, which further validates the practicality and effectiveness of WaDiff. We provide robustness comparisons with other watermarking schemes in the Appendix.
@@ 5.4 Ablation Study
Detection with Different Bit Thresholds
@@
Experiments on Different Sampling Steps
@@
Table 4: The tracing accuracy against two adaptive attacks.
@@
Diff (WaDiff) 77.50% 63.34% 49.42%
Diff (Stega) 0.10% 0.00% 0.00%
Multi-Message (WaDiff) 97.96% 95.66% 91.48%
Multi-Message (Stega) 0.02% 0.00% 0.00%
@@ 5.5 Robustness Analysis with Adaptive Attacks
In addition to the common image augmentations discussed in Section 5.3, we explore two additional countermeasures to evaluate the robustness of our watermarking scheme. For the Diff Attack, we adopt the method from DiffPure [23] on adversarial purification and use a clean diffusion model to purify our watermarked images. For the Multi-Message Attack, we train a surrogate watermark encoder and embed a different watermark message into the watermarked example. All implementation details are provided in the Appendix. We evaluate two adaptive attacks on both WaDiff and StegaStamp and present the experimental results in Table 4. Our results demonstrate that WaDiff exhibits superior robustness compared to StegaStamp in these two adaptive attacks. This can be attributed to the integrated design of our watermarks, which are less susceptible to diffusion denoising and are not easily removed by simply injecting post-hoc information.
@@ 6 Conclusion and Limitations
In this paper, we provide an efficient and robust watermarking framework WaDiff to not only detect whether an image is generated from our model but also identify the specific user who generated the image. With a pre-trained watermark decoder, WaDiff seamlessly incorporates a user-specific watermark as a conditioned input and applies fingerprinting to the generated contents during the image generation process. Extensive experiments demonstrate that WaDiff achieves accurate identification performance among millions of users and remains robust under various data augmentations. While our method still exhibits slightly inferior performance compared to the post-hoc watermark method, however, our WaDiff fingerprints in the generation process which is more stealthy and hard to circumvent in practice. We hope that our work can pioneer the secure auditing of AI-generated content so that we can ensure that model usage aligns with regulatory requirements, compliance, and ethical standards, thereby enhancing the integrity of the generative models.
@@ Appendix 0.A Implementation Details
In this section, we provide additional implementation details of our experiments.
@@ 0.A.4 Consistency Loss for ImageNet Diffusion Model
In this section, we provide more details on the implementation of the consistency loss for the ImageNet diffusion. Instead of directly aligning the output of our WaDiff model with the original output, we take an additional step by applying Equation 4 to both outputs. This gives us the single-step-reverse image $\hat{\mathbf{x}}_{0}^{t}$ for our WaDiff model and $\mathbf{x}_{0}^{t}$ for the original model. To measure the consistency between these images, we adopt the perceptual loss by replacing the consistency loss in Equation 6 with $\mathcal{L}_{c}(b(\hat{\mathbf{x}}_{0}^{t}),b(\mathbf{x}_{0}^{t}))$, where $b$ is a well-trained backbone for feature extraction. Similarly, in our approach, we use the first 23 layers of a pre-trained VGG-16 as $b$ which is the same for training the StageStamp.
@@ 0.B.6 Different Schedulers
We also conducted experiments using different schedulers for image generation, including both the PLMS and DPMSolver. These experiments were performed on the Stable Diffusion model using the official implementation available in the original code repository. For the DPMSolver, we utilized 15 sampling steps due to its high sampling efficiency. Our observations reveal that our method consistently achieves high tracing accuracy across various schedulers. Specifically, with the PLMS scheduler, we obtained a tracing accuracy of 98.20% for Trace $10^{4}$, 96.00% for Trace $10^{5}$, and 93.26% for Trace $10^{6}$. With the DPMSolver scheduler, we achieved a tracing accuracy of 97.98% for Trace $10^{4}$, 95.72% for Trace $10^{5}$, and 92.40% for Trace $10^{6}$. These results highlight the robustness and effectiveness of our method across different scheduling approaches.

### v2

@@ 4.2 Rooting Fingerprints in Diffusion Process
End-to-end Fine-tuning.
Instead of fine-tuning the entire diffusion model, we selectively fine-tune the watermark projector $\mathcal{P}$ and the first input block while keeping the remaining weights unchanged. We adopt this approach for two reasons. First, we observe that fine-tuning the first block is sufficient for effective watermark injection. Our method achieves comparable performance to fine-tuning the entire architecture but with faster speed and lower memory cost. Additionally, we have empirically observed that fine-tuning the entire model can lead to unstable generation, resulting in watermarked images with significantly compromised generative quality. A detailed analysis of fine-tuning various model subsections is provided in the Appendix. Formally, we separate the whole parameters $\theta$ into two sets: $\theta_{head}$, representing the parameters of $\mathcal{P}$ and the first input block, and $\theta_{tail}$, representing the remaining weights, i.e., $\theta=\{\theta_{head},\theta_{tail}\}$. During fine-tuning, only $\theta_{head}$ is optimized while $\theta_{tail}$ is fixed. In sum, we incorporate both optimization objectives above and formulate the fine-tuning process as:
$\min_{\theta_{head}}\mathbb{E}_{\mathbf{x},i,t}[\mathbb{I}(t\leq\tau)(\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t,i}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t}))+\eta\mathcal{L}_{m}(\mathcal{D}(\mathbf{x}_{0,i}^{t}),\mathbf{w}_{i}))$ (7)
$+\mathbb{I}(t>\tau)\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t,null}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t}))],$
@@
Figure 3: Overview of our proposed watermarking framework. The top two figures illustrate our fine-tuning process. For $t>\tau$, we solely focus on preserving image consistency and incorporate a null watermark. For $t\leq\tau$, we integrate the normal watermark and introduce an additional message retrieval loss to embed watermarks. The inference stage is depicted below, where we inject the null watermark when $t>\tau$ and transition it to the payload watermark when $t\leq\tau$.
@@ 4.3 Detection and Identification
Once the diffusion model is fingerprinted, the generated content for each $u_{i}$ will contain specific information that is conditioned by $\mathbf{w}_{i}$. Given a candidate $\mathbf{p}$, we can then recover the source watermark $\mathbf{w}_{s}=\mathcal{D}(\mathbf{p})$ using the pre-trained $\mathcal{D}$ model. Once we have recovered $\mathbf{w}_{s}$, we can utilize it for both detection and identification. To detect whether $\mathbf{p}$ belongs to our model, we set a bit threshold $\tau_{b}$ and calculate the number of matched bits $M_{i}=\mathbf{w}_{s}\odot\mathbf{w}_{i}$, where $\odot$ represents the XNOR function. If $\frac{M_{i}}{n}>\tau_{b}$, we will conclude that $\mathbf{p}$ is generated from our model; otherwise, it is generated from other sources. For identification among $m$ users, we can determine the owner by finding the corresponding user ID whose watermark $\mathbf{w}_{i}$ has the best match with $\mathbf{w}_{s}$, which can be formulated as follows:
$\arg\max_{i}M_{i},\ \ i\in\{1\dots m\}.$ (8)
@@ 5 Experiments
To evaluate the identification performance of our proposed WaDiff, we implement our method on two widely used diffusion models: a text-to-image latent diffusion model and an unconditional diffusion model. We conduct comprehensive comparisons with other watermarking strategies including a traditional strategy DwtDct [6], a deep-learning-based steganography technique StegaStamp [32], which we employed in our pre-training stage, and a recently proposed training-free framework Tree-Ring [34]. Note that we have not included the Stable Signature [11] in our comparison since it considered the model distribution setting which is different from ours. Besides, it necessitates re-training diverse latent decoders, which is not scalable in our identification task. We also conduct thorough ablation studies and provide detailed analysis in the following sections.
5.1 Experimental Settings
Model and Dataset.
For the text-to-image latent diffusion model, we utilize the Stable Diffusion V1.4 as our pre-trained model, and for the unconditional diffusion model, we utilize the 256 $\times$ 256 ImageNet diffusion model. To fine-tune the diffusion models, we randomly select 5000 images with corresponding text descriptions from the training set of MS-COCO 2014 [18] for the Stable Diffusion and 5000 images from the training set of ImageNet [8] for the 256x256 ImageNet diffusion model.
Trace $\text{Trace Acc}=\frac{N_{\text{correct}}}{N_{\text{total}}},$ (9)
Implementation Details.
All experiments were conducted utilizing 8 NVIDIA A100 GPUs. The fine-tuning of both diffusion models involved employing an AdamW optimizer [20] with a learning rate of $1e^{-4}$. For the Stable Diffusion model, we set $\tau$ to 500 and $\eta$ to 0.05, while for the ImageNet diffusion model, we selected $\tau$ as 400 and $\eta$ as 0.25. Notably, in addition to directly aligning the predicted noise of the ImageNet diffusion model with the original model, we empirically found that aligning the single-step-reverse images resulted in improved image quality. Implementation details and additional hyperparameter evaluations are deferred in the Appendix. We fine-tune 40 epochs for the Stable Diffusion model and 25 epochs for the ImageNet diffusion model. For the Stable Diffusion model, we set the default guidance scale to 7.5 and used text descriptions from the validation set of MS-COCO 2014 [18] as prompts. We adopt the watermark length to 48 by default, which is commonly used in previous work [11], and provide experiments of other lengths in the Appendix. We adopted the DDIM [30] sampler with 50 sampling steps as default for both models. More implementation details on training baseline methods and the watermark decoder pre-training is in the Appendix.
@@ 5.2 Detection and Identification Results
In this section, we present the main experimental results of our method. For the detection task, we use 5000 watermarked images along with another 5000 clean images sampled from the original pre-trained diffusion model to calculate the AUC. As for owner identification, we evaluate our method using different sizes of user pools, ranging from ten thousand to one million users. For each user pool, we randomly select 1000 users and generate 5 images per user, resulting in a total of 5000 images. The tracing accuracy is then calculated based on these watermarked images. For SSIM calculation, we first randomly select 200 distinct initial noises and generate a group of 5 images for each noisy vector with different keys. Within each group, the SSIM metric was computed by comparing the similarity between each pair of images. To calculate the FID difference, we evaluate the generative images (for both watermarked and originally generative contents) on the MS-COCO 2014 training set for Stable Diffusion and on the ImageNet [8] training set for the ImageNet diffusion model. We present the experimental results in Table 1. The results demonstrate that our method achieves comparable performance in both detection and identification tasks when compared to the post-hoc StegaStamp, achieving a tracing accuracy of 97.71% and an AUC of 1 on average. We defer further discussion on the comparison with post-hoc methods to the Appendix. In terms of image consistency, our method achieves an average of 0.998 SSIM between pairs of images with different watermarks, which significantly surpasses that of Tree-Ring. Besides, our method only imposes a negligible impact on the original generation quality by a slight increase on the original FID, which is comparable to the post-hoc watermark. Note that the Tree-Ring achieves nearly zero performance on the identification task. This might be attributed to the imprecise latent inversion and the continuous watermarking space, which makes watermarked examples less distinguishable when multiple users are engaged.
@@ 5.3 Watermark Robustness Analysis
To evaluate our watermarking strategy against potential data augmentations, we adopt five commonly used augmentations including image resizing, image blurring, color jitter, Gaussian noising, and JPEG compression. Specifically, for resizing, we randomly resize the width and height of images within a range of 30% to 80% of their original size; for image blurring, we adopt the kernel size to 20; for color jitter, we randomly select from saturation (factor 1.5), contrast (factor 1.5) and sharpness (factor 1.5); for Gaussian noising, we add a Gaussian noise with $\sigma=0.1$; for the JPEG compression, we select the compression quality as 50. We also consider a combinational augmentation that incorporates image resizing, JPEG compression, and color jitter simultaneously. We report the identification performance under various augmentations in Table 2. The experimental results demonstrate the overall robustness of our method against diverse image augmentations, yielding an average tracing accuracy of 88.05% for the Stable Diffusion model and 89.7% for the ImageNet diffusion model. It is noteworthy that our method maintains an average AUC of 0.999 under various data augmentations and achieves identification performance of over 76% among one million users even under combined data augmentation, which further validates the practicality and effectiveness of WaDiff. We provide robustness comparisons with other watermarking schemes in the Appendix.
@@ 5.4 Ablation Study
Detection with Different Bit Thresholds.
@@
Experiments on Different Sampling Steps.
@@
Table 4: The tracing accuracy against two adaptive attacks, where DA and MA are short for the Diff and Multi-Message attack respectively.
@@
DA (WaDiff) 77.50% 63.34% 49.42%
DA (Stega) 0.10% 0.00% 0.00%
MA (WaDiff) 97.96% 95.66% 91.48%
MA (Stega) 0.02% 0.00% 0.00%
@@ 5.5 Robustness Analysis with Adaptive Attacks
In addition to the commonly used image augmentations discussed in Section 5.3, we investigate two additional countermeasures to assess the robustness of our watermarking scheme. For the Diff Attack, we adopt a method similar to DiffPure [24], which uses a pre-trained diffusion model to purify our watermarked images. In our experiments, we set the number of diffusion and denoising steps to 20 and utilize the original DDPM sampler for this attack. In the Multi-Message Attack, we train a surrogate watermark encoder with the same architecture as in our experiments, but using different training data. We then embed a random watermark message into the watermarked example. Evaluation results against two adaptive attacks on both WaDiff and StegaStamp are presented in Table 4. Our results demonstrate that WaDiff exhibits superior robustness compared to StegaStamp in these two adaptive attacks. This can be attributed to the integrated design of our watermarks, which are less vulnerable to diffusion denoising and are not easily removed by directly injecting post-hoc information.
@@ 6 Conclusion and Limitations
In this paper, we provide an efficient and robust watermarking framework WaDiff to not only detect whether an image is generated from our model but also identify the specific user who generated the image. WaDiff seamlessly incorporates a user-specific watermark as a conditioned input and applies fingerprinting to the generated contents during the image generation process. Extensive experiments demonstrate that WaDiff achieves accurate identification performance among a large number of users and remains robust under various data augmentations. While WaDiff still exhibits slightly inferior performance compared to the post-hoc watermarking method, our method fingerprints in the generation process which is more stealthy and hard to circumvent in practice. We hope that our work can pioneer the secure auditing of AI-generated content so that we can ensure that model usage aligns with regulatory requirements, compliance, and ethical standards, thereby enhancing the integrity of the generative models.
@@ Appendix 0.A Implementation Details
In this section, we provide additional implementation details of our experiments. We implement our method based on the official code of the Stable Diffusion [^1] https://github.com/CompVis/stable-diffusion and the ImageNet diffusion model [^2] https://github.com/openai/guided-diffusion.
@@ 0.A.4 Consistency Loss for ImageNet Diffusion Model
In this section, we provide more details on the implementation of the consistency loss for the ImageNet diffusion. Instead of directly aligning the output of our WaDiff model with the original output, we take an additional step by applying Equation 4 to both outputs. This gives us the single-step-reverse image $\mathbf{x}_{0,i}^{t}$ for our WaDiff model and $\mathbf{x}_{0}^{t}$ for the original model. To measure the consistency between these images, we adopt the perceptual loss by replacing the consistency loss in Equation 6 with $\mathcal{L}_{c}(b(\mathbf{x}_{0,i}^{t}),b(\mathbf{x}_{0}^{t}))$, where $b$ is a well-trained backbone for feature extraction. Similarly, in our approach, we use the first 23 layers of a pre-trained VGG-16 as $b$ which is the same for training the StageStamp.
@@ 0.B.6 Different Schedulers
We also conducted experiments using different schedulers for image generation, including both the PLMS and DPMSolver. These experiments were performed on the Stable Diffusion model using the official implementation available in the original code repository. For the DPMSolver, we utilized 15 sampling steps due to its high sampling efficiency. Our observations reveal that our method consistently achieves high tracing accuracy across various schedulers. Specifically, with the PLMS scheduler, we obtained a tracing accuracy of 98.20%, 96.00%, and 93.26% for $10^{4}$, $10^{5}$, and $10^{6}$ users respectively. With the DPMSolver scheduler, we achieved a tracing accuracy of 97.98%, 95.72%, and 92.40% for $10^{4}$, $10^{5}$, and $10^{6}$ users respectively. These results highlight the robustness and effectiveness of our method across different scheduling approaches.